# Roll in the Tanks! Measuring Left-wing Extremism on Reddit at Scale

### v3

@@ Abstract.
Social media’s role in the spread and evolution of extremism is a focus of intense study. Online extremists have been involved in the spread of online hate, mis- and disinformation, and real-world violence. However, most existing work has focuses on right-wing extremism. In this paper, we perform a first of its kind large-scale measurement study exploring left-wing extremism. We focus on “tankies,” a left-wing community that first arose in the 1950s in support of hardline actions of the USSR and has evolved to support what they call “Actually Existing Socialist” countries, e.g., CCP-run China, the USSR, and North Korea. We collect and analyze 1.3M posts from 53K authors from tankie subreddits, and explore the position of tankies within the broader far-left community on Reddit. Among other things, we find that tankies are clearly on the periphery of the larger far-left community. When examining the contents of posts, we find misalignments and conceptual homomorphisms that confirm the description of tankies in the theoretical work. We also discover that tankies focus more on state-level political events rather than social issues. Our findings provide empirical evidence of the distinct positioning and discourse of left-wing extremist groups on social media.
@@ 1. Introduction
At the same time, there has been a steady rise in political rhetoric characterizing mainstream political parties as far-left extremists, scapegoating the far-left for violent activities (e.g., claiming Antifa orchestrated the January 6th Insurrection (Beer, 2021), accusing far-left extremists of planning and organizing violence during protests after George Floyd’s death (Devine, 2020), and blaming left-wing extremists for setting forest fires in Oregon (Harris, 2020)). While most of this rhetoric is easily dismissed as absurd (e.g., President Joe Biden does not fit any definition of far-left that we have found in the literature), many of the characteristics and behavior we associate with right-wing extremism online have historically applied to hardline left-wing extremists as well. For example, spreading mis- and disinformation from unreputable or overtly biased sources (Dutkiewicz and Stecuła, 2022). Despite the political rhetoric, and a history of violence and chaos attributed to far-left extremists, there are essentially no large-scale, data-driven studies of the far-left on social media, let alone far-left extremists. This lack of empirical research leaves a gap in our understanding of whether and how far-left extremism manifests online.
@@
In this paper, we address this gap by conducting the first large-scale measurements of the far-left on social media, using a variety of measurement techniques. We focus primarily on a large left-wing community known as tankies. Historically, tankies were supporters of hard-line Soviet actions (Glastonbury, 1998); more Stalinist than Leninist. The name originates from Soviets using tanks to put down rebellions in eastern Europe (Robinson, 2011). More recently, tankies have grown to support the actions of the CCP in China, a currently operational “Actually Existing Socialist” (AES) country. Notably, their support extends beyond just AES countries, often siding with or excusing anti-NATO, non-socialist, autocratic regimes, including Putin-controlled Russia (Carey, 2022). Regardless of their historical ideology, tankies have recently shown behavior similar to the right-wing extremists (e.g., denying the Uyghur genocide (Roberts, 2020)). Given the parallels in behavior and the lack of measurement studies on left-wing extremism, it is crucial to quantitatively analyze these communities to understand their prevalence and impact on social media platforms.
@@
  • RQ1: What is the scale and growth trajectory of tankie communities on Reddit? As a relatively unknown community in the academic literature, the first step to advancing our understanding is to identify the online tankie community and measure their community growth.
  • RQ2: Where are tankies positioned in the greater left-wing community? While tankies have been characterized in the media and the existing literature as on the extreme end of the left-wing spectrum, at the time of this writing, there is no quantitative evidence to support this positioning.
  • RQ3: What do tankies discuss and how does it differ from other far-left communities? We aim to analyze and measure the tankie discourse, comparing it with other far-left communities to better understand its unique characteristics.
@@
Contributions. In summary, this paper makes several contributions. First, we provide the first large-scale measurements of the far-left on social media and introduce a quantitative framework for discovering and measuring far-left extremists. We identify and quantify the tankie activity on Reddit. We find that tankies had more active users than any other far-left subreddits, their user base has been steadily increasing over time, and that there is a drop in monthly active users in every far-left subreddit except tankies after the Russian invasion of Ukraine (RQ1). Additionally, we measure user migration patterns from other far-left communities to tankies, which coincided with tankies becoming the most popular far-left community on Reddit prior to being banned (RQ1). We then characterize the tankie community on Reddit. To the best of our knowledge, this paper provides the first social media characterization of tankies in the literature across a multitude of disciplines. When looking at tankies’ positioning within our network of ideological subreddits, we find they are indeed at the periphery of the far-left (RQ2), and should thus be considered left-wing extremists. We further show that tankies differ from the rest of the far-left in several striking ways. By developing a metric to measure topic prevalences, we find that tankies give less priority to core left-wing issues like healthcare and housing, instead prioritizing state-level political discussions (RQ3). Finally, we measure linguistic and behavioral differences, finding that tankies show consistent vocabulary misalignments when compared to other far-left communities, exhibit more toxicity, and share more links from revisionist news sources (RQ3).
@@ 2. Background & Related Work
Studies on Extremist Online Communities. Online communities have led to an increased ability for individuals to express their views and connect with others who share similar ideologies. Although the vast majority of these communities are benign, a small subset of them are extremist in nature, espousing views that are well outside the mainstream and often promoting violence (Gaudette et al., 2020; Scrivens et al., 2020). These extremist online communities can exist across the political spectrum, including right-wing and left-wing, as well as religious and other forms of ideologies. Furthermore, it has been observed that there can be similarities and overlap in the user bases of these communities. For example, research by Mame et al. has shown that anti-feminist communities can serve as gateways to the far-right, with significant overlap between the Manosphere and the alt-right observed in their studies on Reddit and YouTube (Mamié et al., 2021). Moreover, while the psycho-political profiles of left and right-wing extremists display considerable diversity (Van Hiel, 2012), these extremist communities may exhibit similarities in their responses to specific events, e.g., left-wing and right-wing extremists fighting against ISIS in Syria (Koch, 2021).
Studies on Islamist extremism investigated the role of social media in the spread of extremist ideologies. Research has shown that these extremists use social media platforms to disseminate propaganda, recruit new members, coordinate attacks, amplify the voices of extremist leaders, and create a sense of community among like-minded individuals (Berger and Morgan, 2015; Piazza and Guler, 2021; Benigni et al., 2017). A comparative study (Jasko et al., 2022) of the use of political violence by left-wing, right-wing, and Islamist extremists found that individuals and attacks associated with left-wing causes are likely to be less violent. Yet, another study (Duran, 2021) analyzing ideologically driven homicides in the US from 1990 to 2020 observed that while fatal violence by right-wing extremists surpasses that of their left-wing counterparts, the intensity of the former’s violent acts has seen a surge post-2015.
@@ 3. Measuring Tankie Activity
Identifying ideology subreddits. To understand tankies’ ideological stance, we initially need to identify prominent ideological subreddits, specifically those leaning left. To accomplish this, we use a list (reddit, 2011) previously applied for such identifications (Shen and Rose, 2021). Subsequently, we carry out a comparative analysis between communities from this list that demonstrate similarity to the tankie ideology. This list has 43 subreddits representing a range of ideologies, from r/Conservative to r/Anarchism, including numerous subvariants, e.g., r/AnarchoPacifism. After manual examination, we remove one non-English subreddit (r/piratenpartei) and those that are defunct or private. Further manual examination shows r/Communist, r/Anarchist, and r/feminisms are not the largest subreddits for their respective ideologies. So, we incorporate the largest subreddits explicitly dedicated to them (r/communism, r/Anarchism, and r/Feminism).
@@
Community Growth. Despite being an old community, tankies have historically exerted minimal tangible influence within the broader context. To understand how things have changed, we first measure the growth and prevalence of the tankies community on Reddit compared to other far-left communities (RQ1) by computing the monthly active user counts (MAU) for each far-left community in our dataset. In the following section (Section 4), we detail our approach to identifying these far-left communities. Subsequently, we measure migration patterns from other far-left communities to tankies to better understand tankies’ growth within far-left ideologies.
Figure 3 shows the MAUs of the largest tankie subreddit, r/GenZedong, compared to non-tankie subreddits. Users active in multiple tankie subreddits are counted only once while calculating tankie MAUs to prevent overestimation. Additional data, including other tankie subreddits and cumulative tankie user counts, are presented in Figure 7 in Appendix B. Starting in July 2020, the tankies community saw a large influx of active users, doubling in about six months and growing larger than the user base of r/communism. As of April 2021, tankies had more active users than any other far-left subreddits. MAU for tankies doubles again between November 2021 and March 2022, reaching 11,872 active users. r/socialism has the second highest MAU of far-left subreddits after tankies with 3,559 (30% of tankies for that month). We also see a decline in monthly active users in every far-left subreddit except tankies starting in February 2022, which is the month Russia invaded Ukraine (John and Murphy, 2022).
To understand the significance of the tankies community’s growth, we run a Mann-Kendall Test (Mann, 1945; Kendall, 1975), a non-parametric test for analyzing monotonic trends in time series data, on our MAU time series. For $\alpha=0.01$, we find tankies are the only community with a statistically significant increase in MAU ($z>0$ and $|z|>z(1-\alpha/2)$, $p<0.01$). We also observe a strong positive correlation between tankies’ MAU and time ( $\rho=0.97$, $p<0.01$) when applying Spearman’s Correlation, indicating that the tankies user base has been steadily increasing over time. These results raise an important question: are these completely new users getting involved in far-left ideologies, or is the user-base of existing far-left subreddits migrating to tankies over time?
@@
Figure 5. The relationship between user bases of ideological subreddits. Far-Left subreddits are colored in red (with non-tankies colored light red), Leftists are colored in gray, Capitalists are colored in green, and Feminists are colored in blue. A stronger red hue indicates a higher similarity between the user-bases of two subreddits.
Takeaways. Tankies on Reddit experienced a notable growth in monthly active users (MAU) since July 2020, surpassing other far-left subreddits in user activity from April 2021 onwards. This indicates that, before facing restrictions on Reddit, tankies became more popular than the other far-left communities we examined (RQ1). Furthermore, while other far-left subreddits experienced a decline in MAU starting in February 2022, the month Russia invaded Ukraine, tankies continued their growth trajectory. This divergence suggests that tankies may have been uniquely impacted by the geopolitical events of that period. One possible explanation for this growth could be the alignment of tankie ideologies with pro-Russian narratives (as we show later in Section 5), which might have attracted users sympathetic to these views during the conflict. Alternatively, the growth could be influenced by coordinated efforts from external actors seeking to amplify certain ideological perspectives, though further investigation would be needed to substantiate such claims. Additionally, their migratory patterns indicate that tankies became less reliant on established far-left communities once they reached a certain level of user base.
@@ 4. Ideological Position of Tankies
In this section, we measure the prominence of tankies within the far-left ideology spectrum and their engagement with other far-left communities (RQ2). To achieve this, we construct a graph of subreddits based on user posting activity, detect communities to identify clusters related to tankie subreddits, and measure user-base similarities between other subreddits in our dataset.
Graph Construction & Community Detection Our first step is to detect communities of similar ideological subreddits. To do this, we follow a methodology similar to that of (Squire, 2019). First, we construct a graph where each node represents a subreddit and the edges between the nodes represent the number of users they have in common. These edges are weighted using Bonacich’s method (Bonacich, 1972), which weighs membership affiliations based on their relative strength within the network. This methodology considers how one’s group membership might be influenced by affiliations to other groups, unveiling structural patterns not evident in unprocessed overlap data.
@@
Far-Left (Red Cluster). March (March, 2008) defines far-left political parties as those that see themselves further left than social democracy. Tankies are indeed clustered with subreddits representing well-known far-left ideologies, e.g., r/communism, r/socialism, r/Anarchism, r/Marxism, and r/IWW (Industrial Workers of the World, a syndicalist labor union (Cole et al., 2017)) representing nearly 9,000 workers across North America (IWW, 2022)). We also see r/alltheleft in this cluster, which has a far-left identity, similar to communist, socialist, and anarchist subreddits (Massoc and Lubda, 2021). This cluster does not split into more clusters when we recursively apply the Louvain algorithm.
@@
Takeaways. Overall, we find that the user-base of tankies is definitively positioned on the periphery of far-left ideologies (RQ2). Our analysis reveals that users within tankie subreddits display stronger intraconnectivity than other subreddits in the broader far-left community, underscoring their unique identity. Furthermore, tankie subreddits exhibit limited cross-engagement with other far-left subreddits, suggesting a degree of ideological isolation. This insularity may reinforce distinct narratives in tankie discourse, deepening their ideological separation from other left-wing groups.

### v4

@@ Abstract.
Social media’s role in the spread and evolution of extremism is a focus of intense study. Online extremists have been involved in the dissemination of online hate, mis- and disinformation, and real-world violence. While the majority of research has focused on right-wing extremism, recent real-world incidents have highlighted the potential for far-left extremists to engage in violence and cause real-world harm as well. In this paper, we present the first large-scale measurement of left-wing extremism on social media. Analyzing 1.3 million posts from 53,000 authors from tankie subreddits, we focus on “tankies,” a left-wing community that first arose in the 1950s in support of hardline actions of the USSR and has evolved to support what they call “Actually Existing Socialist” countries, e.g., CCP-run China, the USSR, and North Korea. Among other things, our analysis reveals that these groups occupy the periphery of the broader far-left community on Reddit, and their discourse distinctively focus on state-level politics and support for authoritarian regimes, rather than on social justice issues. Finally, we show that tankies have high toxicity scores and use pejorative language, mirroring toxicity patterns reported for other online extremist communities. Our findings provide empirical evidence of the distinct positioning and discourse of left-wing extremist groups on social media.
copyright: nonejournalyear: 2026doi: XXXXXXX.XXXXXXXjournal: TSC
@@ 1. Introduction
At the same time, there has been a steady rise in political rhetoric characterizing mainstream political parties as far-left extremists, scapegoating the far-left for violent activities. Examples include claims that antifascist activists orchestrated the January 6th Insurrection (Beer, 2021), allegations that left-wing extremists were responsible for violence during the George Floyd protests (Devine, 2020), or even accusations of left-wing groups setting forest fires in Oregon (Harris, 2020). Much of this rhetoric can be dismissed as political posturing, given that mainstream liberal or left-of-center figures do not fit any reasonable definition of far-left extremism (e.g., the 46th President of the United States, Joe Biden, does not fit any definition of far-left that we have found in the literature).
However, real-world events demonstrate that authoritarian-oriented communities on the far-left also present tangible risks that go beyond political rhetoric. For example, in May 2025, two Israeli embassy staff members were killed outside the Capital Jewish Museum in Washington, D.C., in an attack that investigators linked to far-left extremists (Yousef, 2025). In another example, a son of a CIA deputy director was influenced by pro-authoritarian and anti-Western narratives circulating online, ultimately joining Russian forces in Ukraine, where he was killed in combat (De Luce, 2025). These cases highlight that online spaces associated with far-left extremists are not just a theoretical concern, but can also be a source of real-world recruitment and violence.
Figure 1. Conceptual overview of how tankies differ from other far-left communities on Reddit.
Despite such incidents, and a history of violence and chaos attributed to far-left extremists, there are essentially no large-scale, data-driven studies of the far-left on social media, let alone far-left extremists. This lack of empirical research leaves a gap in our understanding of whether and how far-left extremism manifests online.
@@
In this paper, we address this gap by conducting the first large-scale measurements of the far-left on social media, using a variety of measurement techniques. Following March (March, 2008), we use far-left to refer to actors that place themselves to the left of social democracy and advocate a fundamental transformation of capitalist societies. We focus primarily on a large left-wing community known as tankies. In our analysis, we treat tankies as one specific current within this broader far-left ecosystem. We compare their networks and discourse to other prominent far-left communities to identify when, and in what respects, they adopt more extreme positions and behaviors. Historically, tankies were supporters of hard-line Soviet actions (Glastonbury, 1998); more Stalinist than Leninist. The name originates from Soviets using tanks to put down rebellions in eastern Europe (Robinson, 2011). More recently, tankies have grown to support the actions of the CCP in China, a currently operational “Actually Existing Socialist” (AES) country. Notably, their support extends beyond just AES countries, often siding with or excusing anti-NATO, non-socialist, autocratic regimes, including Putin-controlled Russia (Carey, 2022). Regardless of their historical ideology, tankies have recently shown behavior similar to the right-wing extremists (e.g., denying the Uyghur genocide (Roberts, 2020)). Given the parallels in behavior and the lack of measurement studies on left-wing extremism, it is crucial to quantitatively analyze these communities to understand their prevalence and impact on social media platforms.
@@
  • RQ1: What is the scale and growth trajectory of tankie communities on Reddit? As a relatively unknown community in the academic literature, the first step to advancing our understanding is to identify the online tankie community and measure their community growth. Establishing how large, active, and fast-growing these communities are is a necessary baseline for assessing whether they are a marginal niche or have a substantial extremist presence on Reddit.
  • RQ2: Where are tankies positioned in the greater left-wing community? While tankies have been characterized in the media and the existing literature as on the extreme end of the left-wing spectrum, at the time of this writing, there is no quantitative evidence to support this positioning. By locating tankies within a network of ideological subreddits on Reddit built from shared user participation, we can determine whether they occupy a peripheral position that reflects ideological isolation from the rest of the far-left. Showing that tankies are structurally isolated in this way provides quantitative support for treating them as a distinct left-wing extremist community rather than simply another far-left subreddit.
  • RQ3: What do tankies discuss and how does it differ from other far-left communities? We aim to analyze and measure the tankie discourse, comparing it with other far-left communities to better understand its unique characteristics. This allows us to determine whether tankies simply intensify themes common to the far-left or instead prioritize distinct issues and actors associated with authoritarian, geopolitically aligned extremism.
Figure 1 depicts a conceptual overview for tankies based on our main empirical findings and compares tankies to other far-left communities. It highlighs how they differ in their relative scale and ideological network position, their focus on themes and narratives, and their discourse and language characteristics.
@@
Contributions. In summary, this paper makes several contributions. First, we provide the first large-scale measurements of the far-left on social media and introduce a quantitative framework for discovering and measuring far-left extremists. We identify and quantify the tankie activity on Reddit. We find that tankies had more active users than any other far-left subreddits, their user base has been steadily increasing over time, and that there is a drop in monthly active users in every far-left subreddit except tankies after the Russian invasion of Ukraine (RQ1). Additionally, we measure user migration patterns from other far-left communities to tankies, which coincided with tankies becoming the most popular far-left community on Reddit prior to being banned (RQ1). We then characterize the tankie community on Reddit. To the best of our knowledge, this paper provides the first social media characterization of tankies in the literature across a multitude of disciplines. When looking at tankies’ positioning within our network of ideological subreddits, we find they are indeed at the periphery of the far-left (RQ2), and should thus be considered left-wing extremists. We further show that tankies differ from the rest of the far-left in several striking ways. By developing a metric to measure topic prevalence, we find that tankies give less priority to core left-wing issues like healthcare and housing, instead prioritizing state-level political discussions (RQ3). Finally, we measure linguistic and behavioral differences, finding that tankies show consistent vocabulary misalignments when compared to other far-left communities, and share more links from revisionist news sources (RQ3). We also find that tankies exhibit high toxicity and use pejorative language, in line with prior work on other online extremist communities (Hine et al., 2017; Zannettou et al., 2018; Grover and Mark, 2019) (RQ3). By situating tankies within the broader far-left, our study shows how authoritarian and geopolitically aligned extremists emerges and differentiates itself from other left-wing communities, despite sharing many ideological reference points with them.
@@ 3. Measuring Tankie Activity
Identifying ideology subreddits. To understand tankies’ ideological stance, we initially need to identify prominent ideological subreddits, specifically those leaning left. To accomplish this, we use a list (reddit, 2011) previously applied for such identifications (Shen and Rose, 2021). Subsequently, we carry out a comparative analysis between communities from this list that demonstrate similarities to the tankie ideology. This list has 43 subreddits representing a range of ideologies, from r/Conservative to r/Anarchism, including numerous subvariants, e.g., r/AnarchoPacifism. After manual examination, we remove one non-English subreddit (r/piratenpartei) and those that are defunct or private. Further manual examination shows r/Communist, r/Anarchist, and r/feminisms are not the largest subreddits for their respective ideologies. So, we incorporate the largest subreddits explicitly dedicated to them (r/communism, r/Anarchism, and r/Feminism).
@@
Community Growth. Despite being an old community, tankies have historically exerted minimal tangible influence within the broader context. To understand how things have changed, we first measure the growth and prevalence of the tankies community on Reddit compared to other far-left communities (RQ1) by computing the monthly active user counts (MAU) for each far-left community in our dataset. Specifically, for each far-left subreddit and month, we compute the monthly active user counts (MAU), defined as the number of distinct accounts that post at least once in that subreddit during that month. We then compare the MAU trajectory of tankies to that of other far-left subreddits, allowing us to assess whethere tankies remain a marginal niche or have become one of the main hubs of far-left. In the following section (Section 4), we detail our approach to identifying these far-left communities. Subsequently, we measure migration patterns from other far-left communities to tankies to better understand tankies’ growth within far-left ideologies.
Figure 4 shows the MAUs of the largest tankie subreddit, r/GenZedong, compared to non-tankie subreddits. Additional data, including other tankie subreddits and cumulative tankie user counts, are presented in Figure 9 in Appendix B. Starting in July 2020, the tankies community saw a large influx of active users, doubling in about six months and growing larger than the user base of r/communism. As of April 2021, tankies had more active users than any other far-left subreddits. MAU for tankies doubles again between November 2021 and March 2022, reaching 11,872 active users. r/socialism has the second highest MAU of far-left subreddits after tankies with 3,559 (30% of tankies for that month). We also see a decline in monthly active users in every far-left subreddit except tankies starting in February 2022, which is the month Russia invaded Ukraine (John and Murphy, 2022).
To understand the significance of the tankies community’s growth, we run a Mann-Kendall Test (Mann, 1945; Kendall, 1975), a non-parametric test for analyzing monotonic trends in time series data, on our MAU time series. For $\alpha=0.01$, we find tankies are the only community with a statistically significant increase in MAU ($z>0$ and $|z|>z(1-\alpha/2)$, $p<0.01$). We also observe a strong positive correlation between tankies’ MAU and time ( $\rho=0.97$, $p<0.01$) when applying Spearman’s Correlation, indicating that the tankies user base has been steadily increasing over time. These results raise an important question: are these completely new users getting involved in far-left ideologies, or is the user base of existing far-left subreddits migrating to tankies over time?
@@
Takeaways. Tankies on Reddit experienced a notable growth in monthly active users (MAU) since July 2020, surpassing other far-left subreddits in user activity from April 2021 onwards. This indicates that, before facing restrictions on Reddit, tankies became more popular than the other far-left communities we examined (RQ1). Furthermore, while other far-left subreddits experienced a decline in MAU starting in February 2022, the month Russia invaded Ukraine, tankies continued their growth trajectory. This divergence suggests that tankies may have been uniquely impacted by the geopolitical events of that period. One possible explanation for this growth could be the alignment of tankie ideologies with pro-Russian narratives (as we show later in Section 5), which might have attracted users sympathetic to these views during the conflict. Alternatively, the growth could be influenced by coordinated efforts from external actors seeking to amplify certain ideological perspectives, though further investigation would be needed to substantiate such claims. Additionally, their migratory patterns indicate that tankies became less reliant on established far-left communities once they reached a certain level of user base. Taken together, these growth and migration patterns show that tankies evolved from a relatively marginal subculture into a rapidly expanding hub within the far-left. This justifies our subsequent focus on their position and discourse, since the distinctive behaviors we document in later sections concern a community whose audience is actively expanding rather than a static fringe.
@@
Figure 6. The relationship between user bases of ideological subreddits. Far-Left subreddits are colored in red (with non-tankies colored light red), Leftists are colored in gray, Capitalists are colored in green, and Feminists are colored in blue. A stronger red hue indicates a higher similarity between the user bases of two subreddits.
@@ 4. Ideological Position of Tankies
In this section, we measure the prominence of tankies within the far-left ideology spectrum and their engagement with other far-left communities (RQ2). To achieve this, we construct a graph of subreddits based on user posting activity, detect communities to identify clusters related to tankie subreddits, and measure user base similarities between other subreddits in our dataset.
Graph Construction & Community Detection Our first step is to detect communities of similar ideological subreddits. To do this, we follow a methodology similar to that of (Squire, 2019). First, we construct a graph where each node represents a subreddit and the edges between the nodes represent the number of users they have in common. These edges are weighted using Bonacich’s method (Bonacich, 1972), which normalizes co-membership based on the full $2\times 2$ contingency table of shared and non-shared members across each pair of subreddits. This controls for differences in community size and produces a normalized coefficient in $[0,1]$ that reflects the strength of membership overlap between two communities. This methodology allows us to distinguish meaningful structural ties from overlaps that arise simply because some subreddits are much larger than others, unveiling structural patterns not evident in unprocessed overlap data.
@@
Far-Left (Red Cluster). Building on March’s definition of the far-left (March, 2008), we label this cluster as far-left, since tankies are indeed clustered with subreddits representing well-known far-left ideologies, e.g., r/communism, r/socialism, r/Anarchism, r/Marxism, and r/IWW (Industrial Workers of the World, a syndicalist labor union (Cole et al., 2017)), representing nearly 9,000 workers across North America (IWW, 2022). We also see r/alltheleft in this cluster, which has a far-left identity, similar to communist, socialist, and anarchist subreddits (Massoc and Lubda, 2021). This cluster does not split into more clusters when we recursively apply the Louvain algorithm.
@@
Takeaways. Overall, we find that the user base of tankies is definitively positioned on the periphery of far-left ideologies (RQ2). Our analysis reveals that users within tankie subreddits display stronger intraconnectivity than other subreddits in the broader far-left community, underscoring their unique identity. Furthermore, tankie subreddits exhibit limited cross-engagement with other far-left subreddits, suggesting a degree of ideological isolation. This insularity may reinforce distinct narratives in tankie discourse, deepening their ideological separation from other left-wing groups.